OpenStudy (kamille):

Hello, I need to solve this problem If John works alone, he will do the work per 4hours. If Mark works alone, he'll do the work per 3h. At first John and Mark worked together and done 7/8 of all work. How much time it would tike for John if he does all the work left (1/8)? Any tips?

OpenStudy (kamille):

hey, @hartnn I can try to rewrite the problem, if you dont understand what I am asking for

hartnn (hartnn):

can you first find, in how much time they will complete the entire work, working together ?

OpenStudy (kamille):

\[\frac{ x }{ 3 }+\frac{ x }{4 }=1\]?

OpenStudy (kamille):

actually, i dont know how to solve these type of problems:/

hartnn (hartnn):

ok, if A can do the work in 'x' time and B can do the same work in 'y' time, then they both together can do the same work in xy/(x+y) time. understand/remember this.

OpenStudy (kamille):

oh,okay so if John does work in x time, Mark does in y time, it means: \[\frac{ 3*4 }{ 3+4 }=\frac{ 12 }{ 7 }\]

hartnn (hartnn):

yes, they complete the work in 12/7 hours totally, so in what time will they complete 7/8 th of the work ?

OpenStudy (kamille):

is it a proportion? I will do something like this